# A reconstructed discontinuous approximation for distributed elliptic control problems

### v1

@@ Abstract.
In this paper, we present and analyze an internal penalty discontinuous Galerkin method for the distributed elliptic optimal control problems. It is based on a reconstructed discontinuous approximation which admits arbitrarily high-order approximation space with only one unknown per element. Applying this method, we develop a proper discretization scheme that approximates the state and adjoint variables in the approximation space. Our main contributions are twofold: (1) the derivation of both a priori and a posteriori error estimates of the $L^{2}$-norm and the energy norms, and (2) the implementation of an efficiently solvable discrete system, which is solved via a linearly convergent projected gradient descent method. Numerical experiments are provided to verify the convergence order in a priori estimate and the efficiency of a posteriori error estimate.
@@ 2.1. Model Problem
Let $\Omega_{u}\subset\mathbb{R}^{m}$ and $\Omega\subset\mathbb{R}^{d}(m,d=2,3)$ be convex polygonal or polyhedral domains. Take the control space $U=L^{2}(\Omega_{u}),$ the state space $V=H^{2}(\Omega)$ and the admissible set $U_{ad}\subset U$ which is closed and convex. Consider the following optimal control problem:
@@
In addition, we assume $f\in L^{2}(\Omega),\phi\in H^{\frac{3}{2}}(\partial\Omega),B:L^{2}(\Omega_{u})\rightarrow L^{2}(\Omega)$ is a bounded linear operator.
The state equation of problem (2.1) can be divided into two independent parts as
Notice that $y_{2}$ does not rely on $u$, we denote ${g_{1}}(y_{1}):=g(y_{1}+y_{2})$. Define a bilinear form $a:H_{0}^{1}(\Omega)\times H_{0}^{1}(\Omega)\rightarrow\mathbb{R}$ as
$a(y_{1},w):=(A\nabla y_{1},\nabla w)_{L^{2}(\Omega)}.$
Denote the optimal solution to (2.1) by $(y^{\ast},u^{\ast})$, then we have the state equation
(2.2) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y^{\ast})=f+Bu^{*}\quad\mathrm{in}\ \Omega,\\ &y^{\ast}=\phi\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
and the relation $y_{1}^{\ast}=y^{\ast}-y_{2}$. Define the adjoint variable $p^{\ast}$ such that
Consider the Lagrange function
$L(y_{1}(u),u):={g_{1}}(y_{1}(u))+j(u)-a(y_{1}(u),p^{\ast})+(f+Bu,p^{\ast})\quad\forall u\in U_{ad}.$
Here we use (,)($\cdot\,$$\cdot\,$) to denote the $L^{2}$ inner-product $(\,\cdot\,,\,\cdot\,)_{L^{2}}$. Notice that $L(y_{1}(u),u)=g_{1}(y_{1}(u))+j(u),\forall u\in U_{ad},$ we have
$L(y_{1}(u^{\ast}+t(v-u^{\ast})),u^{\ast}+t(v-u^{\ast}))\geq L(y_{1}^{\ast},u^{\ast}),\quad\forall v\in U_{ad},\ t\in[0,1].$
Differentiate the above inequality at $t=0$ in the direction $v-u^{*}.$ We have
$(g_{1}^{\prime}(y_{1}^{*})y_{1}^{\prime}(u^{*})+j^{\prime}(u^{*}))(v-u^{*})-a(y_{1}^{\prime}(u^{*})(v-u^{*}),p^{*})+(B(v-u^{*}),p^{*})\geq 0.$
Since
we take $w=y_{1}^{\prime}(u^{*})(v-u^{*})\in H_{0}^{1}(\Omega)$ and derive
$(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.$
@@
Assumption 2.
The functionals $j(u)$ and $g(y)$ can be written in the form
which means that
$(j^{\prime}(u),v)=\int_{\Omega_{u}}\tilde{j}^{\prime}(u)v\mathrm{d}x,\quad(g^{\prime}(y),w)=\int_{\Omega}\tilde{g}^{\prime}(y)w\mathrm{d}x,$
for $\forall u,v\in U,y,w\in Y.$
[…]
@@ 2.3. The Reconstructed Discontinuous Approximation Method
Here $\mathcal{E}_{K}=\{e\in\mathcal{E}_{h}\mid e\subset\partial K\}.$ So it does not change under uniform refinement of $\mathcal{T}_{h}.$
@@
Let $(y_{h}^{m*},u_{h}^{*})$ be the solution to (2.13). The Lagrange function
$L(y_{h}^{m}(u_{h}),u_{h}):={g}(y_{h}^{m}(u_{h}))+j(u_{h})-(g^{\prime}(y_{h}^{m*}),y_{h}^{m}(u_{h})-F_{h}(f+Bu_{h},\phi)).$
@@
We have $L(y_{h}^{m}(u_{h}),u_{h})=g(y_{h}^{m}(u_{h}))+j(u_{h}),\forall u_{h}\in U_{adh}.$ So
$L(y_{h}^{m}(u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast})),u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast}))\geq L(y_{h}^{m*},u_{h}^{*}),\quad\forall v_{h}\in U_{adh},\ t\in[0,1].$
$(\tilde{g}^{\prime}(y_{h}^{m*})y_{h}^{m\prime}(u_{h}^{*})+\tilde{j}^{\prime}(u_{h}^{*}),v_{h}-u_{h}^{*})$
- $(\tilde{g}^{\prime}(y_{h}^{m*}),y_{h}^{m\prime}(u_{h}^{*})(v_{h}-u_{h}^{*})-F_{h}(B(v_{h}-u_{h}^{*}),0))\geq 0.$
Here we use Assumption 2. Then
$(\tilde{j}^{\prime}(u_{h}^{*}),v_{h}-u_{h}^{*})+(\tilde{g}^{\prime}(y_{h}^{m*}),F_{h}(B(v_{h}-u_{h}^{*}),0))\geq 0.$
For $\forall w\in L^{2}(\Omega),$ we have
@@ 4.1. A General Result
$-\nabla\cdot(A\nabla Sy)=f+Bu_{h}\quad Sy|_{\partial\Omega}=\phi,$
@@ 5. Numerical Results
Next we show the results of $e_{K_{u}}$ for $m=1$. Consistent with Example 4, we plot $\min\{e_{K_{u}},4\}$ to avoid the influence of extreme values in $e_{K_{u}}$. In this case, our restriction is $U_{ad}=\{u\in L^{2}(\Omega_{u})\mid 0\leq u\leq 1\ \mathrm{a.e.\ in}\ \Omega_{u}\}$. Two clear borderlines are evident in the results, corresponding to the conditions $\{u_{d}-s(p^{*})=0\}$ and $\{u_{d}-s(p^{*})=1\}$. This is illustrated in Figures 17 and 19. Outside these boundaries, $e_{K_{u}}$ remains under 4.

### v2

@@ Abstract.
In this paper, we present and analyze an interior penalty discontinuous Galerkin method for the distributed elliptic optimal control problems. It is based on a reconstructed discontinuous approximation which admits arbitrarily high-order approximation space with only one unknown per element. Applying this method, we develop a proper discretization scheme that approximates the state and adjoint variables in the approximation space. Our main contributions are twofold: (1) the derivation of both a priori and a posteriori error estimates of the $L^{2}$-norm and the energy norms, and (2) the implementation of an efficiently solvable discrete system, which is solved via a linearly convergent projected gradient descent method. Numerical experiments are provided to verify the convergence order in a priori error estimate and the efficiency of a posteriori error estimate.
@@ 2.1. Model Problem
Let $\Omega_{u},\Omega\subset\mathbb{R}^{d}(d=2,3)$ be convex polygonal or polyhedral domains. Take the control space $U=L^{2}(\Omega_{u}),$ the state space $V=H^{2}(\Omega)$ and the admissible set $U_{ad}\subset U$ which is closed and convex. Consider the following optimal control problem:
@@
In addition, we assume $f\in L^{2}(\Omega),\phi\in H^{\frac{3}{2}}(\partial\Omega),B:L^{2}(\Omega_{u})\rightarrow L^{2}(\Omega)$ is a bounded linear operator. We further assume that
Assumption 1.
The functionals $j(u)$ and $g(y)$ can be written in the form
which means that
$j^{\prime}(u)(v)=\int_{\Omega_{u}}\tilde{j}^{\prime}(u)v\mathrm{d}x,\quad g^{\prime}(y)(w)=\int_{\Omega}\tilde{g}^{\prime}(y)w\mathrm{d}x,$
for $\forall u,v\in U,y,w\in Y.$
[…]
@@
Define a bilinear form $a:H_{0}^{1}(\Omega)\times H_{0}^{1}(\Omega)\rightarrow\mathbb{R}$ as
$a(y,w):=(A\nabla y,\nabla w)_{L^{2}(\Omega)}.$
To derive the optimality condition of problem (2.1), we separate the state equation into two independent parts as
Notice that $y_{2}$ does not rely on $u$, we denote ${g_{1}}(y_{1}(u)):=g(y_{1}(u)+y_{2})$. Denote the optimal solution to (2.1) by $(y^{\ast},u^{\ast})$, then
(2.2) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y^{\ast})=f+Bu^{*}\quad\mathrm{in}\ \Omega,\\ &y^{\ast}=\phi\quad\mathrm{on}\ \partial\Omega,\end{aligned}\right$.
and $y_{1}^{\ast}=y^{\ast}-y_{2}$. Define the adjoint variable $p^{\ast}$ such that
Define the cost functional $J(y_{1}(u),u):={g_{1}}(y_{1}(u))+j(u),$ we have
$J(y_{1}(u^{\ast}+t(v-u^{\ast})),u^{\ast}+t(v-u^{\ast}))\geq J(y_{1}^{\ast},u^{\ast}),\quad\forall v\in U_{ad},\ t\in[0,1].$
Differentiate the above inequality at $t=0$ in the direction $v-u^{*},$ then
$(g_{1}^{\prime}(y_{1}^{*})y_{1}^{\prime}(u^{*})+j^{\prime}(u^{*}))(v-u^{*})\geq 0,\quad\forall v\in U_{ad}.$
Notice that $a(y_{1}(v),p^{*})=(f+Bv,p^{*})$ for $\forall v\in U_{ad}$, we have
$a(y_{1}^{\prime}(u^{*})(v-u^{*}),p^{*})-(B(v-u^{*}),p^{*})=0,\quad\forall v\in U_{ad}.$
Here we use (,)($\cdot\,$$\cdot\,$) to denote the $L^{2}$ inner-product $(\,\cdot\,,\,\cdot\,)_{L^{2}}$. Since
we take $w=y_{1}^{\prime}(u^{*})(v-u^{*})\in H_{0}^{1}(\Omega)$ and derive
$(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.$
@@ 2.3. The Reconstructed Discontinuous Approximation Method
Here $\mathcal{E}_{K}=\{e\in\mathcal{E}_{h}\mid e\subset\partial K\}.$ So it is invariant under uniform refinement of $\mathcal{T}_{h}.$
@@
Let $(y_{h}^{m*},u_{h}^{*})$ be the solution to (2.13). Notice that
$J(y_{h}^{m}(u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast})),u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast}))\geq J(y_{h}^{m*},u_{h}^{*}),\quad\forall v_{h}\in U_{adh},\ t\in[0,1].$
$(\tilde{g}^{\prime}(y_{h}^{m*})y_{h}^{m\prime}(u_{h}^{*})+\tilde{j}^{\prime}(u_{h}^{*}),v_{h}-u_{h}^{*})\geq 0,\quad\forall v_{h}\in U_{adh}.$
@@
Using the equality
$(\tilde{g}^{\prime}(y_{h}^{m*}),y_{h}^{m}(v_{h})-F_{h}(f+Bv_{h},\phi))=0,\quad\forall v_{h}\in U_{adh},$
we deduce that
$(\tilde{g}^{\prime}(y_{h}^{m*}),y_{h}^{m\prime}(u_{h}^{*})(v_{h}-u_{h}^{*})-F_{h}(B(v_{h}-u_{h}^{*}),0))=0,\quad\forall v_{h}\in U_{adh}.$
In addition, for $\forall w\in L^{2}(\Omega),$
@@
Take $w=B(v_{h}-u_{h}^{*})$ to derive
$(\tilde{j}^{\prime}(u_{h}^{*})+B^{*}F_{h}(\tilde{g}^{\prime}(y_{h}^{m*}),0),v_{h}-u_{h}^{*})\geq 0,\quad\forall v_{h}\in U_{adh}.$
@@ 4.1. A General Result
$-\nabla\cdot(A\nabla Sy)=f+Bu_{h},\quad Sy|_{\partial\Omega}=\phi,$
@@ 5. Numerical Results
Next we show the results of $e_{K_{u}}$ for $m=1$. Consistent with Example 4, we plot $\min\{e_{K_{u}},4\}$ to avoid the influence of extreme values in $e_{K_{u}}$. In this case, our restriction is $U_{ad}=\{u\in L^{2}(\Omega_{u})\mid 0\leq u\leq 1\ \mathrm{a.e.\ in}\ \Omega_{u}\}$. Two borderlines are evident in the results, corresponding to the conditions $\{u_{d}-s(p^{*})=0\}$ and $\{u_{d}-s(p^{*})=1\}$. This is illustrated in Figures 17 and 19. Outside these boundaries, $e_{K_{u}}$ remains under 4.